# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

### v2

@@ 5.1 Predictability of Stock Return Using Earnings-Price Ratio
In spite of a prominent economist and statistician with a national reputation, Irving Fisher suffered devastating personal finance losses during the 1929 Great Depression (Dimand and Geanakoplos,, 2005). Were he alive today, he would have mixed feelings about the fact that return predictability remains inconclusive after a century of extensively studies. Among the various predictors examined, the predictive power of valuation ratios, particularly the earnings-price ratio, has been subject to widespread discussions. In this section, we investigate the predictability of stock returns using the log earnings-price ratio while incorporating high dimensional mixed-root control covariates.
@@ 5.1.1 Data
Our analysis focuses on predicting the monthly return of the S&P 500 index, calculated as $\text{Return}_{t}=\log(P_{t})-\log(P_{t-1}),$ where $P_{t}$ refers to S&P 500 (S&P’s Common Stock Price Index: Composite). The primary predictor of interest is the log earnings-price ratio. We obtain it from the variable S&P PE ratio (S&P’s Composite Common Stock: Price-to-Earnings Ratio), denoted as $\text{PE}_{t}$. The log earnings-price ratio is calculated by inverting the original price-earning ratio as $\text{logEP}_{t}=\log(1/\text{PE}_{t}).$
@@
Figure B.1 displays the time series plot of the monthly return of the S&P 500 index and the log earnings-price ratio from January 1960 to April 2024. The log earnings-price ratio exhibits persistent patterns, while the S&P 500 return appears stationary. We further report the AR(1) coefficient estimates and Augmented Dickey–Fuller (ADF) test $p$-values of both series under different sample periods in Table B.1. The S&P 500 return is evidenced to be stationary under the full sample with an ADF test $p$-value below 1%, rejecting the null hypothesis of nonstationarity. Conversely, the log earnings-price ratio shows high persistence with an AR(1) coefficient estimate equal to $0.993$. Given the $p$-value of $0.081$, nonstationarity is not rejected under the 5% significance level. Note that the nonstationary log earnings-price ratio is likely to have predictive power for the stationary monthly return of S&P 500, since we allow for a local-to-zero coefficient to balance the different scales between a stationary outcome and a nonstationary regressor (Phillips,, 2015). Theorem 2 and the paragraph that follows illuminate that inference by XDlasso has the power to detect a wide range of local-to-zero alternatives.
@@
Return predictability can be time-varying and sporadic (Tu and Xie,, 2023). Campbell and Yogo, (2006) discover that the return predictability test results can vary depending on the inclusion of the samples after 1994, and the break is revisited by Zhu et al., (2014). Following this empirical finding, we divide the sample into two periods: pre-1994 (January 1960 - December 1993) and post-1994 (January 1994 - April 2024). The S&P 500 return demonstrates stationarity across both subperiods with ADF $p$-values below 1%. In contrast, the persistence of the log earnings-price ratio is ambiguous. Nonstationarity is evident in the Pre-1994 period with a large ADF test $p$-value $0.776$, while the null hypothesis of nonstationarity is rejected at the 1% significance level in the Post-1994 period. The ambiguity in stationarity of the log earnings-price ratio motivates the use of the XDlasso procedure, a unified approach for both stationary and nonstationary regressors without prior knowledge of their persistence.
@@
Our analysis incorporates high dimensional covariates including all other 110 macroeconomic variables from the FRED-MD dataset. These variables are mixed with stationary and nonstationary time series. To facilitate the stationarization, FRED assigns a transformation code (TCODE) to each variable denoting the recommended transformation of the potentially nonstationary time series into a stationary one. Given the stationary pattern of the dependent variable, in our analysis, we follow the TCODE to transform the 110 time series and subsequently use the transformed variables as covariates. Nevertheless, we highlight that these elementary transformations are not a silver bullet in getting rid of nonstationarity. We perform the ADF test for each of the transformed variables for different sample periods. A nontrivial proportion (9.1%) of the transformed variables still demonstrate nonstationarity based on the ADF test at 5% significance level for both Pre-1994 and Post 1994 subperiods. The high persistence of the log earnings-price ratio and other covariates suggests the necessity of XDlasso.
@@ 5.1.2 Results
We study the one-month-ahead regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes a high dimensional vector that collects all control variables. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the log earnings-price ratio in forecasting the S&P 500 monthly return. In addition to XDlasso and Dlasso, we conduct the IVX inference in a simple regression setup, omitting all control covariates, to demonstrate the necessity of incorporating high dimensional control covariates in practice. We compare the results of XDlasso to those of Dlasso for the high dimensional regression and IVX for a simple regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+u_{t}.$
Table 3: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ : Full sample and subperiods
Sample Periods IVX Dlasso XDlasso
Est. S.E. Est. S.E. Est. S.E.
Full Sample (Jan. 1960 - Apr. 2024) -0.023** 0.010 0.008* 0.006 0.002 0.016
Pre-1994 (Jan. 1960 - Dec. 1993) -0.058** 0.026 0.025** 0.010 0.057* 0.032
Post-1994 (Jan 1994 - Apr. 2024) -0.022 0.015 0.000 0.007 -0.001 0.019
@@
Notes: ”Est.” and ”S.E.” represent the estimates and standard error, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. The tuning parameter for LASSO estimation is selected through 10-fold cross-validation. In XDlasso, instruments are generated based on (2.7) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
@@
Table 3 presents the point estimates and testing results for $\theta_{1}^{*}$. IVX delivers negative point estimates across all sample periods and detects significant effects in the full sample period and Pre-1994 subperiod at the 5% significance level. However, the resulting negative relation between stock return and earning performance contradicts the economic mechanism between the two variables. Obviously, a large number of potential confounding variables are present in the dataset, which may be the culprit in producing the counterintuitive result from the simple regression.
Dlasso, which incorporates the high dimensional controls, reverses the signs of estimates. It returns a positively significant the coefficient for the full sample at the 10% significance level, and detects stronger evidence of predictability, at 5% significance level, in the Pre-1994 period. Further accounting for the potential nonstationarity in the predictors, XDlasso provides different estimates and standard errors from Dlasso. Notably, the standard errors of Dlasso are 60% to 68% smaller than those of XDlasso across sample periods, suggesting Dlasso may underestimate the uncertainty. Consequently, XDlasso reveals that there is no significant predictive power of log earnings-price ratio for stock return in the full sample. Moreover, the evidence of predictor power in Pre-1994 is weakened to 10% level compared with that of Dlasso at 5% level. Our empirical findings are consistent with the general recognition that there is little predictability in the financial market, thereby supporting the efficient market hypothesis.
@@ 5.2 Predictability of Inflation Using Unemployment Rate
It is essential for monetary policymakers to understand the relationship between unemployment and inflation. As Engemann, (2020) pointed out, “The Federal Reserve has a dual mandate to promote maximum sustainable employment and price stability.” First alluded by Fisher, (1926, 1973) though, Phillips, (1958) popularized the Phillips curve — a negative relationship between the level of unemployment and the change rate of money wage rates. There has been a prolonged debate about whether the unemployment is a credible barometer for inflation among not only modern economic studies, but also policymakers. [^4] Mary Daly, San Francisco Fed President, delivered at Daly, (2019) a negative view on the Phillips curve that “As for the Phillips curve… most arguments today center around whether it’s dead or just gravely ill. Either way, the relationship between unemployment and inflation has become very difficult to spot.” John Williams, New York Fed President, expressed a different opinion that “The Phillips curve is the connective tissue between the Federal Reserve’s dual mandate goals of maximum employment and price stability. Despite regular declarations of its demise, the Phillips curve has endured. It is useful, both as an empirical basis for forecasting and for monetary policy analysis.” See Engemann, (2020) for more details. Empirical findings find that inflation rate can be either positively or negatively correlated to unemployment, depending on the shocks to the economies, the policies, and the lag orders (Niskanen,, 2002; Gordon,, 2011, 2013). Given the ongoing debate, We revisit the Phillips curve in a predictive regression framework utilizing the FRED-MD dataset.
@@ 5.2.1 Data
Similar to our first empirical application in Section 5.1, we follow Benati, (2015) to delineate three subperiods in addition to the full sample: Pre-Volcker (January 1960–July 1979), Volcker and Greenspan (August 1979–January 2006), and Bernanke, Yellen, and Powell (February 2006 - April 2024). These subperiods correspond to different eras in U.S. monetary policy, each named after the Federal Reserve chairperson who presided during that time. [^5] Pre-Volcker (before August 1979): This period was characterized by high and volatile inflation, with the Federal Reserve lacking a clear nominal anchor. Volcker and Greenspan (August 1979–January 2006): Volcker, who served as the Federal Reserve Chairman from August 1979 to August 1987, implemented aggressive anti-inflation measures, notably raising interest rates to historically high levels. Greenspan succeeded Volcker and continued to focus on maintaining price stability during his tenure, which contributed to a period of low and stable inflation known as the “Great Moderation”. Bernanke, Yellen, and Powell (February 2006–December 2019): Bernanke’s tenure as Fed Chairman was marked by the Great Recession and the implementation of unconventional monetary policies, such as quantitative easing, aimed at stimulating the economy and preventing deflation. Yellen and Powell carried on these policies. This periodization allows us to examine how the relationship between inflation and unemployment may have evolved across different policy regimes and economic conditions.
@@
Figure B.2 plots the inflation and unemployment rates over our sample period. Visual inspection suggests that the unemployment rate is more persistent than the inflation rate. Table B.2 further reports the AR(1) coefficient estimates and ADF test $p$-values of the inflation and the unemployment rate under different sample periods. The inflation rate demonstrates stationary patterns in most periods, with AR(1) coefficient estimates ranging from 0.54 to 0.64. However, the inflation rate during the pre-Volcker period shows a slight upward trend, and nonstationarity is indicated by the ADF test with a $p$-value of 0.22. In contrast, the inflation rate is found to be stationary in the other two subperiods. The unemployment rate, on the other hand, appears nonstationary with AR(1) coefficient estimates close to 1. The ADF test rejects the null hypothesis of nonstationarity at a 10% significance level for both the full sample and the Volcker-Greenspan period. However, during the Pre-Volcker and Bernanke-Yellen-Powell periods, there is strong evidence of nonstationarity in the unemployment rate. The stationarity of both inflation and the unemployment rate in different periods is again ambiguous, which prompts the use of XDlasso from an agnostic perspective.
@@
In addition to the unemployment rate, we incorporate all other 110 macroeconomic variables from the FRED-MD dataset as controls after the TCODE transformation as in Section 5.1, in view of the pattern of the dependent variable. Still, a significant proportion, from 11% to 17% across subperiods, of the transformed variables exhibit nonstationarity according to the ADF test at 5% significance level.

### v3

@@ 5.1 Predictability of Stock Return Using Earnings-Price Ratio
In financial economics, testing the predictive power of valuation ratios, particularly the earnings-price ratio, on stock returns has been subject to widespread discussions. Much of the literature tests the predictability using univariate predictive regressions (e.g., welch2008comprehensive, zhu2014predictive, and goyal2024comprehensive), but inference can be sensitive to model misspecification arising from omitted variables. Controlling for high dimensional covariates is therefore necessary not only to enhance out-of-sample prediction but also to mitigate the omitted variables for credible and accurate inference. The literature on predictive regression for stock returns has focused on identifying which variables possess significant predictive power for future returns. For better out-of-sample prediction, recent literature has documented gains in forecasting performance from incorporating high-dimensional covariates into predictive regressions; see, for instance, smeekes2018macroeconomic, gu2020empirical and medeiros2021forecasting. The increasing popularity of statistical learning methods with high dimensional predictors calls for suitable inference methods beyond univariate predictive regressions. The goal is to provide investors with signals to adjust their portfolios based on changes in these predictive variables. In this section, we investigate the predictability of stock returns using the log earnings-price ratio in a data-rich environment with high dimensional mixed-root control variables.
@@ 5.1.1 Data
Our analysis focuses on predicting the monthly return of the S&P 500 index, calculated as $\text{Return}_{t}=\log(P_{t})-\log(P_{t-1}),$ where $P_{t}$ refers to S&P 500 (S&P’s Common Stock Price Index: Composite). The primary predictor of interest is the log earnings-price ratio. We obtain it from the variable S&P PE ratio (S&P’s Composite Common Stock: Price-to-Earnings Ratio), denoted as $\text{PE}_{t}$. The log earnings-price ratio is calculated by inverting the original price-earnings ratio as $\text{logEP}_{t}=\log(1/\text{PE}_{t}).$
@@
Figure 4 displays the time series plot of the monthly return of the S&P 500 index and the log earnings-price ratio from January 1960 to April 2025. The log earnings-price ratio exhibits persistent patterns, while the S&P 500 return appears stationary. We further report the AR(1) coefficient estimates and Augmented Dickey–Fuller (ADF) test $p$-values of both series under different sample periods in Table 4. The S&P 500 return is evidenced to be stationary under the full sample with an ADF test $p$-value below 1%, rejecting the null hypothesis of nonstationarity. Conversely, the log earnings-price ratio shows high persistence with an AR(1) coefficient estimate equal to $0.993$. Given the $p$-value of $0.074$, nonstationarity is not rejected at the 5% significance level. Note that the nonstationary log earnings-price ratio can predict the stationary monthly return of S&P 500 in our model, since we allow for a local-to-zero coefficient to balance the different scales between a stationary outcome and a nonstationary regressor (phillips2015halbert). Theorem 2 and the paragraph that follows illuminate that inference by XDlasso has the power to detect a wide range of local-to-zero alternatives.
Figure 4: S&P 500 monthly Return and Log Earnings-price Ratio
Table 4: Persistence of S&P 500 Monthly Return and Log Earnings-Price Ratio
Sample Period S&P 500 Monthly Return Log Earnings-Price Ratio
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2025) 0.227 <<0.01 0.993 0.074
Pre-1994 (Jan. 1960 - Dec. 1993) 0.255 <<0.01 0.994 0.776
Post-1994 (Jan. 1994 - Apr. 2025) 0.200 <<0.01 0.979 <<0.01
Notes: The lag order for the ADF test is set to $\lfloor n^{1/3}\rfloor$ where $n$ is the effective sample size. The exact start and end dates of subperiods are provided in the first column of the table.
@@
Return predictability can be time-varying and sporadic (tu2023penetrating). campbell2006efficient discover that the return predictability test results can vary depending on the inclusion of the samples after 1994, and the break is revisited by zhu2014predictive. Following this empirical finding, we divide the sample into two periods: pre-1994 (January 1960 - December 1993) and post-1994 (January 1994 - April 2025). The S&P 500 return demonstrates stationarity across both subperiods with ADF $p$-values below 1%. In contrast, the persistence of the log earnings-price ratio depends on the sample period considered. Nonstationarity is evident in the Pre-1994 period with a large ADF test $p$-value $0.776$, while the null hypothesis of nonstationarity is rejected at the 1% significance level in the Post-1994 period. The ambiguity in stationarity of the log earnings-price ratio motivates the use of the XDlasso procedure, a unified approach for both stationary and nonstationary regressors without prior knowledge of their persistence.
@@
Besides the monthly return of the S&P 500 index and the log earnings-price ratio, our analysis incorporates high dimensional covariates including all other 110 macroeconomic variables from the FRED-MD dataset. These variables comprise a mixture of stationary and nonstationary time series. In practice, it is common for empirical analysts to transform potentially nonstationary time series into stationary ones to avoid challenges arising from nonstationarity. To facilitate such stationarization, FRED assigns a transformation code (TCODE) to each variable denoting the recommended transformation.
In our analysis, we follow the common practice of using the TCODE to transform the 110 time series, and subsequently use the transformed variables as covariates. Nevertheless, we highlight that these elementary transformations are not a silver bullet in taming nonstationarity. We perform the ADF test for each of the transformed variables for different sample periods. A nontrivial proportion (9.1%) of the transformed variables still demonstrate nonstationarity based on the ADF test at the 5% significance level for both Pre-1994 and Post-1994 subperiods. The high persistence of the log earnings-price ratio and other covariates suggests the necessity of XDlasso.
As highlighted by smeekes2020unit, the predictive performance in regressions using FRED-MD data is sensitive to the transformations. To assess the robustness of our results, we also conduct our analysis using the original (untransformed) FRED-MD time series as covariates. To reduce the impact of highly nonstationary series, we exclude I(2) variables that require second differencing for stationarity according to their TCODE classification. [^4] Appendix C.1 further investigates robustness by considering: (i) excluding nonstationary variables based on integrated orders determined by the bootstrap sequential testing procedure of smeekes2015bootstrap, as reported in smeekes2020unit, and (ii) applying only logarithmic transformations, as indicated by TCODE, without differencing.
@@ 5.1.2 Results
We study the one-month-ahead regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes a high dimensional vector that collects all control variables. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the log earnings-price ratio in forecasting the S&P 500 monthly return. Given the stationary pattern of the dependent variable, we additionally consider specifications where $W_{-1,t-1}$ includes the lagged dependent variable $\text{Return}_{t-1}$. For each model, we apply the wild bootstrapped automatic variance ratio test (kim2009automatic) to the Slasso residuals as a diagnostic test for the martingale difference sequence (m.d.s.) condition in Assumption 1.555This practice serves as a heuristic diagnostic. Demonstrated by our simulation results in Appendix B.4, the variance ratio test on the Slasso residuals $\widehat{u}_{t}$ tends to severely over-reject the m.d.s. condition for $u_{t}$. To the best of our knowledge, at present the literature has no valid testing procedure yet for m.d.s. in high dimensional predictive regressions. It is an open question for future research.
@@
In addition to XDlasso and Dlasso, we conduct the IVX inference in a simple regression setup, omitting all control variables, to demonstrate the necessity of incorporating high dimensional control variables in practice. We compare the results of XDlasso to those of Dlasso for the high dimensional regression and IVX for a simple regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+u_{t}.$
Table 5: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ across sample periods and specifications in stock return prediction
(a) TCODE Transformed Data
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period IVX Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.023** 0.009 0.003 0.074 0.009 0.005 0.216
(Jan. 1960 - Apr. 2025) (0.010) (0.006) (0.015) (0.006) (0.015)
Pre-1994 -0.059** 0.025** 0.059* 0.227 0.024** 0.062* 0.296
(Jan. 1960 - Dec. 1993) (0.026) (0.010) (0.033) (0.010) (0.032)
Post-1994 -0.022 0.002 -0.001 0.053 0.002 -0.001 0.049
(Jan. 1994 - Apr. 2025) (0.015) (0.007) (0.017) (0.007) (0.017)
(b) Untransformed Data: Excluding I(2) Variables Based on TCODE
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period IVX Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.023** 0.013 -0.008 0.001 0.019 0.012 0.811
(Jan. 1960 - Apr. 2025) (0.010) (0.014) (0.011) (0.014) (0.010)
Pre-1994 -0.059** 0.064** -0.312 0.046 0.055* 0.096 0.467
(Jan. 1960 - Dec. 1993) (0.026) (0.031) (0.290) (0.033) (0.070)
Post-1994 -0.022 -0.003 -0.022 0.016 -0.000 -0.004 0.280
(Jan. 1994 - Apr. 2025) (0.015) (0.009) (0.017) (0.008) (0.016)
@@
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The upper and lower panels corresponds to the scenarios where we use TCODE transformed or untransformed time series as covariates, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
@@
Table 5 presents the point estimates and testing results for $\theta_{1}^{*}$. IVX delivers negative point estimates across all sample periods and detects significant effects in the full sample period and Pre-1994 subperiod at the 5% significance level. However, the resulting negative relation between stock return and earning performance contradicts the economic mechanism between the two variables. A large number of potential confounding variables are present in the dataset, which may be the culprit in producing the counterintuitive result from the simple regression.
Dlasso, which accounts for high dimensionality but not nonstationarity, generally reverses the sign of the estimates compared to the simple regression. For the full sample, Dlasso consistently yields positive but statistically insignificant coefficients. Zooming into the Pre-1994 period, Dlasso detects strong evidence of predictability with transformed data, with coefficients significant at the 5% level. This result persists with the untransformed data, though the significance weakens to the 10% level when the lagged dependent variable is included. Accounting for the potential nonstationarity in the predictors, XDlasso provides notably different estimates and standard errors from Dlasso. XDlasso reveals no significant predictive power of log earnings-price ratio for stock return in the full sample and the Post-1994 period. Moreover, for the Pre-1994 period, the evidence of predictability is weakened to the 10% significance level with transformed data, and becomes entirely insignificant with untransformed data. Our empirical findings with XDlasso are consistent with the general recognition that there is little predictability in the financial market, thereby supporting the efficient market hypothesis. The divergence between Dlasso and XDlasso underscores our theoretical prediction: XDlasso mitigates the Stambaugh bias arising from highly persistent regressors while Dlasso does not. Despite the short confidence intervals, Dlasso results can be misleading when nonstationary time series is present, as clearly shown by the illustrative simulation in Section 2.4.
The variance ratio tests on the Slasso residuals provide a diagnostic check on the m.d.s. condition required by our asymptotic theory. The results exhibit stark differences across data specifications: with untransformed data and without the lagged dependent variable, the tests strongly reject the m.d.s. condition ($p$-values ranging from 0.001 to 0.046), suggesting potential model misspecification. Including the lagged return or using transformed data substantially improves the diagnostic results, with $p$-values generally exceeding 0.05. This pattern suggests that both data transformation and including dynamics help satisfy the underlying assumptions, thereby boosting credibility for the XDlasso results under these specifications.
@@ 5.2 Predictability of Inflation Using Unemployment Rate
It is essential for monetary policymakers to understand the relationship between unemployment and inflation. As engemann2020phillips pointed out, “The Federal Reserve has a dual mandate to promote maximum sustainable employment and price stability.” First alluded to by fisher1926statistical; fisher1973discovered though, phillips1958relation popularized the Phillips curve — a negative relationship between the level of unemployment and the change rate of money wage rates. There has been a prolonged debate about whether the unemployment is a credible barometer for inflation among not only modern economic studies, but also policymakers. [^6] Mary Daly, San Francisco Fed President, delivered at daly2019new a negative view on the Phillips curve that “As for the Phillips curve… most arguments today center around whether it’s dead or just gravely ill. Either way, the relationship between unemployment and inflation has become very difficult to spot.” John Williams, New York Fed President, expressed a different opinion that “The Phillips curve is the connective tissue between the Federal Reserve’s dual mandate goals of maximum employment and price stability. Despite regular declarations of its demise, the Phillips curve has endured. It is useful, both as an empirical basis for forecasting and for monetary policy analysis.” See engemann2020phillips for more details. Empirical findings suggest that inflation rate can be either positively or negatively correlated with unemployment, depending on the shocks to the economies, the policies, and the lag orders (niskanen2002death; gordon2011history; gordon2013phillips). Given the ongoing debate, we revisit the Phillips curve in a predictive regression framework utilizing the FRED-MD dataset.
@@ 5.2.1 Data
Similar to our first empirical application in Section 5.1, we follow benati2015long to delineate three subperiods in addition to the full sample: Pre-Volcker (January 1960 - July 1979), Volcker and Greenspan (August 1979 - January 2006), and Bernanke, Yellen, and Powell (February 2006 - April 2025). These subperiods correspond to different eras in U.S. monetary policy, each named after the Federal Reserve chairperson who presided during that time. [^7] Pre-Volcker (before August 1979): This period was characterized by high and volatile inflation, with the Federal Reserve lacking a clear nominal anchor. Volcker and Greenspan (August 1979–January 2006): Volcker, who served as the Federal Reserve Chairman from August 1979 to August 1987, implemented aggressive anti-inflation measures, notably raising interest rates to historically high levels. Greenspan succeeded Volcker and continued to focus on maintaining price stability during his tenure, which contributed to a period of low and stable inflation known as the “Great Moderation”. Bernanke, Yellen, and Powell (February 2006–December 2019): Bernanke’s tenure as Fed Chairman was marked by the Great Recession and the implementation of unconventional monetary policies, such as quantitative easing, aimed at stimulating the economy and preventing deflation. Yellen and Powell carried on these policies. This periodization allows us to examine how the relationship between inflation and unemployment may have evolved across different policy regimes and economic conditions.
@@
Figure 5 plots the inflation and unemployment rates over our sample period. Visual inspection suggests that the unemployment rate is more persistent than the inflation rate. Table 7 further reports the AR(1) coefficient estimates and ADF test $p$-values of the inflation and the unemployment rate under different sample periods. The inflation rate appears stationary in most periods, with AR(1) coefficient estimates ranging from 0.54 to 0.64. However, the inflation rate during the pre-Volcker period shows a slight upward trend, and nonstationarity is indicated by the ADF test with a $p$-value of 0.22. In contrast, the inflation rate is found to be stationary in the other two subperiods. The unemployment rate, on the other hand, appears highly persistent with AR(1) coefficient estimates close to 1. The ADF test rejects the null hypothesis of nonstationarity at a 10% significance level for both the full sample and the Volcker-Greenspan period. However, during the Pre-Volcker and Bernanke-Yellen-Powell periods, there is strong evidence of nonstationarity in the unemployment rate. The stationarity of both inflation and the unemployment rate in different periods is again ambiguous, which prompts the use of XDlasso from an agnostic perspective.
Figure 5: Inflation and Unemployment Rate
Table 7: Persistence of Inflation Rate and Unemployment Rate
Sample Period Inflation Rate Unemployment Rate
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2025) 0.620 0.0183 0.999 0.074
Pre-Volcker (Jan. 1960 - July 1979) 0.640 0.220 0.986 0.406
Volcker-Greenspan (Aug. 1979 - Jan. 2006) 0.623 <<0.01 0.992 0.071
Bernanke-Yellen-Powell (Feb. 2006 - Apr. 2025) 0.537 <<0.01 0.939 0.325
Notes: The lag order for the ADF test is set to $\lfloor n^{1/3}\rfloor$ where $n$ is the effective sample size. The exact start and end dates of subperiods are provided in the first column of the table.
@@
In addition to the unemployment rate, we incorporate all other 110 macroeconomic variables from the FRED-MD dataset as controls after the TCODE transformation. Still, a significant proportion, from 11% to 17% across subperiods, of the transformed variables exhibit nonstationarity according to the ADF test at the 5% significance level. Consistent with Section 5.1, we also perform the analysis using the original (untransformed) FRED-MD time series as covariates. To further address potential concerns regarding the high correlation between control variables and the unemployment rate, we additionally present results excluding variables from the labor market group, as classified by FRED-MD.